OpenStudy (anonymous):

(2x^(2))^(1/3)*(2x)^(1/2)

OpenStudy (anonymous):

sorry express using radicals

OpenStudy (anonymous):

2x^2/3 *2x^1/2 \[\sqrt{2x^3} * \sqrt{2x}\]

OpenStudy (anonymous):

I got that 2/3 by multiplying the powers : 2 *1/3=2/3